5. If the function y = y(x) is such that x dy dx + 2y = x cos x, use the integrating-factor method to show that y = sin x + 2 cos x /x − 2 sin x/x^2 + c/x^2 , where c is an arbitrary constant


1
Expert's answer
2022-02-04T07:55:14-0500

Solution:

"\begin{aligned}

&\text { Given } x \frac{d y}{d x}+2 y=x \cos x \\

&\Rightarrow\left(x \frac{d y}{d x}+2 y\right) \times \frac{1}{x}=x \cos x \times \frac{1}{x} \\

&\Rightarrow \frac{d y}{d x}+\frac{2 y}{x}=\cos x \\

&\Rightarrow \frac{d y}{d x}+\left(\frac{2}{x}\right) y=\cos x

\end{aligned}"

This is a first order linear differential equation of the form

"\frac{d y}{d x}+P y=Q"  Here, "P=\frac{2}{x}\ and\ Q=\cos x"

The integrating factor (I.F) of this differential equation is,

"\begin{aligned}

&\text { I. F }=\mathrm{e}^{\int \mathrm{Pdx}} \\

&\Rightarrow \text { I.F }=\mathrm{e}^{\int \frac{2}{\mathrm{x}} \mathrm{dx}} \\

&\Rightarrow \text { I.F }=\mathrm{e}^{2 \int \frac{1}{\mathrm{x}} \mathrm{dx}}

\end{aligned}

\\


\begin{aligned}

&\text { We have } \int \frac{1}{\mathrm{x}} \mathrm{dx}=\log \mathrm{x}+\mathrm{c} \\

&\Rightarrow \text { I.F }=\mathrm{e}^{2 \log \mathrm{x}}

\end{aligned}"

"\begin{aligned}

&\Rightarrow \mathrm{I} . \mathrm{F}=\mathrm{e}^{\log \mathrm{x}^{2}} \\

{ } \\

{\left[\because \mathrm{m} \log \mathrm{a}=\log \mathrm{a}^{\mathrm{m}}\right]} \\

{\therefore \mathrm{I} . \mathrm{F}=\mathrm{x}^{2}\left[\because \mathrm{e}^{\log \mathrm{x}}=\mathrm{x}\right]}

\end{aligned}"

Hence, the solution of the differential equation is,

"\begin{aligned}

y(I . F)=& \int(Q \times I . F) d x+c \\

\Rightarrow y\left(x^{2}\right)=& \int\left(\cos x \times x^{2}\right) d x+c \\

\Rightarrow y x^{2}=& \int x^{2} \cos x d x+c \\

\Rightarrow y x^{2}=& \int\left(x^{2}\right) \times(\cos x) d x+c \\

\text { Recall } \int f(x) g(x)=f(x)\left[\int g(x) d x\right] \\

&-\int\left[f^{\prime}(x)\left(\int g(x) d x\right)\right] d x+c \\

\Rightarrow y x^{2}=& x^{2}\left[\int \cos x d x\right] \\

&-\int\left[\frac{d}{d x}\left(x^{2}\right)\left(\int \cos x d x\right)\right] d x+c

\end{aligned}"

"\begin{aligned}
\Rightarrow y x^{2} &=x^{2} \sin x-2 \int x \sin x d x+c \\
\Rightarrow y x^{2} &=x^{2} \sin x-2\left\{x\left[\int \sin x d x\right]\right.\\
&\left.-\int\left[\frac{d}{d x}(x)\left(\int \sin x d x\right)\right] d x\right\}+c \\
\Rightarrow y x^{2} &=x^{2} \sin x-2\{x[-\cos x]\\
\Rightarrow y x^{2} &=x^{2} \sin x-2\{-x \cos x\\
\Rightarrow y x^{2} &=x^{2} \sin x \\
&-2\{-x \cos x+\sin x\}+c \\
\Rightarrow y x^{2} &=x^{2} \sin x+2 x \cos x \\
 & \quad-2 \sin x+c \\

\end{aligned}"

 "\mathrm{y}=\sin \mathrm{x}+\frac{2}{\mathrm{x}} \cos \mathrm{x}-\frac{2}{\mathrm{x}^{2}} \sin \mathrm{x}+\frac{\mathrm{c}}{\mathrm{x}^{2}}"

Thus, the solution of the given differential equation is

"y=\sin x+\frac{2}{x} \cos x-\frac{2}{x^{2}} \sin x+\frac{c}{x^{2}}"